OpenStudy (anonymous):

7th grade question Marlin's marble jar had 36 red marbles and 64 blue marbles. What was the ratio of red marbles to blue marbles in the jar?

OpenStudy (thatonegirl_):

your ratio in words will look like this- red:blue. So you have 36 red, and 64 blue. 36:64. But you can simplify this, correct?

OpenStudy (anonymous):

So it's C or D am I righT?

OpenStudy (anonymous):

36 to 64

OpenStudy (anonymous):

A. 1 : 2 B. 9 : 16 C. 2 : 3 D. 3 : 4

OpenStudy (anonymous):

forgot to post the possible answers, please forgive me

OpenStudy (thatonegirl_):

It's going to be B.

OpenStudy (thatonegirl_):

You can divide both of them by 2, 36/2=18. 64/2=32. Now the ratio is 18:32. You can then again divide these by 2. 18/2=9 32/2=16. The ratio is 9:16. Do you get it?

OpenStudy (anonymous):

yeah i do thanks

OpenStudy (thatonegirl_):

No problem